# Xiaoyi Pang1, Zicong Hong2, Ruikai Shi1, Hao Huang1, Jie Zhang1, Song Guo1†

### v1

@@
[…]
HALO: A Unified Vision-Language-Action Model for Embodied Multimodal Chain-of-Thought Reasoning Quanxin Shou1∗, Fangqi Zhu1∗, Shawn Chen1, Puxin Yan3, Zhengyang Yan1, Yikun Miao1,
Xiaoyi Pang1, Zicong Hong2, Ruikai Shi1, Hao Huang1, Jie Zhang1, Song Guo1†
1The Hong Kong University of Science and Technology, 2EPFL, 3Sun Yat-sen University
Correspondence to: qshou@connect.ust.hk, songguo@ust.hk
@@ 2 Related Work
Predominant VLA models are built on top of Vision-Language Models (VLMs) pre-trained on large-scale web data, inheriting rich multimodal knowledge about the world. Representative architectures such as OpenVLA (O’Neill et al., 2024) and FAST (Pertsch et al., 2025) directly extend pre-trained VLMs by introducing discrete action tokens and predicting actions autoregressively to solve embodied tasks. The $\pi_{0}$ series (Black et al., 2024; Intelligence et al., 2025) further introduce a separate action expert along with the pre-trained VLM for flow-based action prediction, achieving better performance and efficiency.
Beyond equipping VLM with only action prediction capability, unified VLA architectures (Bi et al., 2025; Wang et al., 2025; Cen et al., 2025; Lv et al., 2025) have been proposed to enhance dynamic awareness by jointly modeling textual understanding, visual generation, and action prediction. These methods incorporate image or video generation components to predict future environmental states, using either unified (Cen et al., 2025; Wang et al., 2025) or mixed (Bi et al., 2025; Lv et al., 2025) architectures that are trained together with action prediction to capture underlying environment dynamics. However, these methods lack explicit textual or visual reasoning capability, limiting their performance in complex and dynamic environments.
@@
Reasoning is a crucial step towards robotic intelligence (LeCun, 2022). Motivated by the success of Chain-of-Thought reasoning in language models, recent work extends explicit reasoning mechanisms to embodied decision making. MoTVLA (Bi et al., 2025) proposes a textual reasoning framework that performs planning and reasoning prior to action execution. UP-VLA (Zhang et al., 2025), CoTVLA (Zhao et al., 2025) and InternVLA-A1 (Cai et al., 2026) further introduce visual reasoning by generating goal images and videos to guide downstream action prediction.
More recently, ManualVLA (Gu et al., 2025) tries to combine visual and textual reasoning through a shared planning expert that produces both textual instructions and visual subgoals. However, this design couples image and text generation within a single autoregressive model, which may degrade the reasoning capability of the underlying VLM. In contrast, our proposed embodied multimodal chain-of-thought (EM-CoT) employs a Mixture-of-Transformers (MoT) architecture for joint textual reasoning, visual generation, and action prediction. Our decoupled design enables more effective reasoning across visual and textual modalities before executing actions.
@@ 3.2 Unified Architecture
To facilitate training stability, we initialize each expert using the Qwen2.5-1.5B LLM (Yang et al., 2025), yielding a cumulative parameter count of approximately 4.5B. To project heterogeneous data into a unified representation space, we employ modality-specific encoders. For text, we adopt the standard Qwen2.5 tokenizer (Bai et al., 2023), expanding the vocabulary with special control tokens—specifically $\langle\text{think\_start}\rangle$, $\langle\text{vision\_start}\rangle$, and $\langle\text{action\_start}\rangle$, among others—to strictly demarcate the textual reasoning, visual generation, and acting phases, thereby enforcing a structured and interpretable workflow. In the vision domain, we decouple the encoding mechanisms for semantic understanding and visual generation to address their distinct functional requirements. For the understanding task, we employ a ViT (Dosovitskiy, 2020) initialized with SigLIP2 (Tschannen et al., 2025) to leverage robust semantic features. To accommodate diverse real-world inputs without geometric distortion, this encoder is further enhanced with NaViT (Dehghani et al., 2023), enabling the processing of images in their native aspect ratios. For the visual generation task, we utilize a pre-trained VAE (Kingma and Welling, 2013; Labs et al., 2025) optimized for high-fidelity image compression and reconstruction. Finally, for the action modality, given its low-dimensional and physically structured nature, a linear projection suffices to map actions into the model’s hidden dimension while preserving intrinsic structural information. See Appendix A for further architectural details.
@@ 3.4 Training Recipe
  • VQA (Mutilmodal understanding): We utilize LLaVA-NeXT-779k (Liu et al., 2024a) to provide high-quality multimodal conversational priors. This task aligns linguistic instructions with visual contexts via a standard cross-entropy loss, denoted as $\mathcal{L}_{\text{CE}}$.
@@ 4.3 Ablation Study
Effectiveness of the versatile pre-training. As depicted in Panel A, the full versatile pre-training utilizing all three diverse datasets demonstrates unparalleled effectiveness, consistently achieving the highest performance metrics. Conversely, even the selective removal of a single data modality, such as visual generation data (w/o V), leads to a significant degradation in performance, particularly on hard tasks (over 50% reduction). Further progressive exclusion of textual (w/o V+T) and action prediction data (w/o V+T+A) results in a continuously worsening performance trajectory. Notably, without any pre-training (w/o V+T+A), the model’s performance falls to a complete 0% on hard tasks, demonstrating that pre-training is an absolutely foundational requirement for establishing the core competencies necessary to tackle complex problems. Collectively, these findings underscore the indispensable and complementary roles of the versatile pre-training step.
Effectiveness of EM-CoT. As shown in Panel B, both visual subgoal images and textual reasoning are crucial components for EM-CoT, removing either component degrades performance. Combining both reasoning modalities in Halo achieves the highest accuracy, particularly on hard tasks. These results validate the effectiveness of the proposed EM-CoT formulation.
@@ 4.4 Qualitative Results of EM-CoT
To qualitatively assess the efficacy of EM-CoT, we visualize representative rollouts from both the clean and randomized settings of RoboTwin 2.0, as illustrated in Figure 6.
In the clean setting, EM-CoT demonstrates high fidelity, with predicted textual reasoning and visual subgoals closely aligning with the ground truth. The result indicates that Halo effectively masters the underlying task logic, providing a clear, self-generated multimodal roadmap that anchors the final action in a semantically and visually consistent manner.
Under aggressive environmental randomization, EM-CoT still exhibits significant robustness while maintaining performance despite substantial visual shifts. As shown in the last row of Figure 6, Halo correctly identifies targets and plans trajectories even when object appearances and backgrounds significantly differ from the training distribution. This suggests our EM-CoT paradigm allows the agent to perform true semantic reasoning rather than relying on simple pattern matching. By enforcing a deliberate reasoning process, the model effectively performs low-level control guided by high-level reasoning under the challenging out-of-distribution scenarios.
@@ A.2 Multimodal Encoders
  • Visual Input: We employ separate visual encoders to handle visual understanding and generation independently. For visual understanding, we utilize a ViT (Dosovitskiy, 2020) encoder initialized with SigLIP2-so400m/14 (Tschannen et al., 2025) (pre-trained at 384×384). To enhance high-resolution perception, we interpolate position embeddings to support inputs up to 980×980 and integrate NaViT (Dehghani et al., 2023) to process images in their native aspect ratios. A two-layer MLP projector is adopted to align visual features with LLM’s embedding space. For visual generation, we leverage the pre-trained VAE (Kingma and Welling, 2013) from FLUX (Labs et al., 2025), featuring a downsample ratio of 8 and a latent channel of 16. The VAE latents are processed by flattening 2×2 spatial patches into vectors, which are then projected by a linear layer to match the hidden dimension of the LLM. Conversely, a VAE decoder is employed similarly to reconstruct images from latents. All VAE parameters remain frozen during training.

### v2

# arXiv:2602.21157v2[cs.RO] 27 Feb 2026

@@
Abstract
@@ 2 Related Work
Predominant VLA models are built on top of Vision-Language Models (VLMs) pre-trained on large-scale web data, inheriting rich multimodal knowledge about the world. Representative architectures such as OpenVLA (O’Neill et al., 2024) and FAST (Pertsch et al., 2025) directly extend pre-trained VLMs by introducing discrete action tokens and predicting actions autoregressively to solve embodied tasks. The $\pi_{0}$ series (Black et al., 2024; Intelligence et al., 2025) further introduce a separate action expert along with the pre-trained VLM for flow-based action prediction, achieving better performance and efficiency. Beyond equipping VLM with only action prediction capability, unified VLA architectures (Bi et al., 2025; Wang et al., 2025; Cen et al., 2025; Lv et al., 2025) have been proposed to enhance dynamic awareness by jointly modeling textual understanding, visual generation, and action prediction. These methods incorporate image or video generation components (Zhu et al., 2025, 2026) to predict future environmental states, using either unified (Cen et al., 2025; Wang et al., 2025) or mixed (Bi et al., 2025; Lv et al., 2025) architectures that are trained together with action prediction to capture underlying environment dynamics. However, these methods lack explicit textual or visual reasoning capability, limiting their performance in complex and dynamic environments.
@@
Reasoning is a crucial step towards robotic intelligence (LeCun, 2022). Motivated by the success of Chain-of-Thought reasoning in language models, recent work extends explicit reasoning mechanisms to embodied decision making. MoTVLA (Bi et al., 2025) proposes a textual reasoning framework that performs planning and reasoning prior to action execution. UP-VLA (Zhang et al., 2025), CoTVLA (Zhao et al., 2025) and InternVLA-A1 (Cai et al., 2026) further introduce visual reasoning by generating goal images and videos to guide downstream action prediction. More recently, ManualVLA (Gu et al., 2025) tries to combine visual and textual reasoning through a shared planning expert that produces both textual instructions and visual subgoals. However, this design couples image and text generation within a single autoregressive model, which may degrade the reasoning capability of the underlying VLM. In contrast, our proposed embodied multimodal chain-of-thought (EM-CoT) employs a Mixture-of-Transformers (MoT) architecture for joint textual reasoning, visual generation, and action prediction. Our decoupled design enables more effective reasoning across visual and textual modalities before executing actions.
@@ 3.2 Unified Architecture
To facilitate training stability, we initialize each expert using the Qwen2.5-1.5B LLM (Yang et al., 2025), yielding a cumulative parameter count of approximately 4.5B. To project heterogeneous data into a unified representation space, we employ modality-specific encoders. For text, we adopt the standard Qwen2.5 tokenizer (Bai et al., 2023), expanding the vocabulary with special control tokens—specifically $\langle\text{think\_start}\rangle$, $\langle\text{vision\_start}\rangle$, and $\langle\text{action\_start}\rangle$, among others—to strictly demarcate the textual reasoning, visual generation, and acting phases, thereby enforcing a structured and interpretable workflow. In the vision domain, we decouple the encoding mechanisms for semantic understanding and visual generation to address their distinct functional requirements. For the understanding task, we employ a ViT (Dosovitskiy, 2020) initialized with SigLIP2 (Tschannen et al., 2025) to leverage robust semantic features. To accommodate diverse real-world inputs without geometric distortion, this encoder is further enhanced with NaViT (Dehghani et al., 2023), enabling the processing of images in their native aspect ratios. For the visual generation task, we utilize a pre-trained VAE (Kingma & Welling, 2013; Labs et al., 2025) optimized for high-fidelity image compression and reconstruction. Finally, for the action modality, given its low-dimensional and physically structured nature, a linear projection suffices to map actions into the model’s hidden dimension while preserving intrinsic structural information. See Appendix A for further architectural details.
@@ 3.4 Training Recipe
  • VQA (Mutilmodal understanding): We use LLaVA-NeXT-779k (Liu et al., 2024a) to provide high-quality multimodal conversational priors. This task aligns linguistic instructions with visual contexts via a cross-entropy loss, denoted as $\mathcal{L}_{\text{CE}}$.
@@ 4.3 Ablation Study
Effectiveness of the versatile pre-training
. As depicted in Panel A, the full versatile pre-training utilizing all three diverse datasets demonstrates unparalleled effectiveness, consistently achieving the highest performance metrics. Conversely, even the selective removal of a single data modality, such as visual generation data (w/o V), leads to a significant degradation in performance, particularly on hard tasks (over 50% reduction). Further progressive exclusion of textual (w/o V+T) and action prediction data (w/o V+T+A) results in a continuously worsening performance trajectory. Notably, without any pre-training (w/o V+T+A), the model’s performance falls to a complete 0% on hard tasks, demonstrating that pre-training is an absolutely foundational requirement for establishing the core competencies necessary to tackle complex problems. Collectively, these findings underscore the indispensable and complementary roles of the versatile pre-training step.
Effectiveness of EM-CoT
. As shown in Panel B, both visual subgoal images and textual reasoning are crucial components for EM-CoT, removing either component degrades performance. Combining both reasoning modalities in Halo achieves the highest accuracy, particularly on hard tasks. These results validate the effectiveness of the proposed EM-CoT formulation.
@@ 4.4 Qualitative Results of EM-CoT
To qualitatively assess the efficacy of EM-CoT, we visualize representative rollouts from both the clean and randomized settings of RoboTwin 2.0, as illustrated in Figure 6. In the clean setting, EM-CoT demonstrates high fidelity, with predicted textual reasoning and visual subgoals closely aligning with the ground truth. The result indicates that Halo effectively masters the underlying task logic, providing a clear, self-generated multimodal roadmap that anchors the final action in a semantically and visually consistent manner. Under aggressive environmental randomization, EM-CoT still exhibits significant robustness while maintaining performance despite substantial visual shifts. As shown in the last row of Figure 6, Halo correctly identifies targets and plans trajectories even when object appearances and backgrounds significantly differ from the training distribution. This suggests our EM-CoT paradigm allows the agent to perform true semantic reasoning rather than relying on simple pattern matching. By enforcing a deliberate reasoning process, the model effectively performs low-level control guided by high-level reasoning under the challenging out-of-distribution scenarios.
@@ A.2 Multimodal Encoders
  • Visual Input: We employ separate visual encoders to handle visual understanding and generation independently. For visual understanding, we utilize a ViT (Dosovitskiy, 2020) encoder initialized with SigLIP2-so400m/14 (Tschannen et al., 2025) (pre-trained at 384×384). To enhance high-resolution perception, we interpolate position embeddings to support inputs up to 980×980 and integrate NaViT (Dehghani et al., 2023) to process images in their native aspect ratios. A two-layer MLP projector is adopted to align visual features with LLM’s embedding space. For visual generation, we leverage the pre-trained VAE (Kingma & Welling, 2013) from FLUX (Labs et al., 2025), featuring a downsample ratio of 8 and a latent channel of 16. The VAE latents are processed by flattening 2×2 spatial patches into vectors, which are then projected by a linear layer to match the hidden dimension of the LLM. Conversely, a VAE decoder is employed similarly to reconstruct images from latents. All VAE parameters remain frozen during training.